# How Far Are We From AGI?

### v1

@@ 3.3.1 AI Interface to Other AI agents
  • Agents as teachers and learners. On one hand, stronger AGI models often act as oracles to provide ‘supervision’ to inferior ones, from tuning on data from better models (Taori et al., 2023; Gu et al., 2023) to prompt-engineering-based approaches (Huang et al., 2022a; Jiang et al., 2023a; Fu et al., 2023b), there emerges the concept of model knowledge distillation. In the field of language processing, it is common to use a teacher system to label and expand existing data by directly taking the teacher’s answer (Gilardi et al., 2023; Hsieh et al., 2023; Li et al., 2022a; Sun et al., 2024a; Ding et al., 2023a) or using more advanced techniques such as CoT prompting (Ramnath et al., 2023; Li et al., 2023g), or creating new data for subsequent models to distill useful and compact knowledge from large-scale data (Li et al., 2023b; Javaheripi et al., 2023). Similar paradigms are applied in computer vision and multimodal domains for better model training and deployment. One of the most prevailing methods is utilizing the GPT-4V model to label answers in various tasks (Shu et al., 2023; Liu et al., 2023d; Li et al., 2023j).
    In conventional solutions, a better AGI model severs as the role of the teacher, however, there is a growing trend for less competent AGI models to provide insights for aligning stronger ones with or beyond human perception, known as superalignment (Burns et al., 2023). This has proven to be effective in empowering higher capacity than the teacher model with a vanilla fine-tuning strategy and distilled data from the smaller model. Recent studies have begun to explore how leveraging weaker models to enhance the performance of stronger ones can be applied across various domains. Chen et al. (2024a) introduced the innovative idea of iteratively harnessing the capacity of weaker models to improve the efficacy of more powerful counterparts. (Ji et al., 2024) developed an efficient alignment paradigm that learns the correctional residuals between aligned and unaligned responses from a weaker model. (Sun et al., 2024b) employs a less advanced system, trained on simpler tasks, to guide more capable models for tackling hard reasoning tasks (e.g., level 4-5 MATH problems). In the vision domain, (Guo et al., 2024) presented compelling evidence that weak-to-strong generalization may outperform finetuning methods in certain scenarios. This interaction between different models requires acquiring knowledge from selected models step-by-step, forming the sequential interface.
@@
    Looking beyond multi-modal tasks, more specific and advanced agent systems that can handle more complex computer applications such as web browsing (Mialon et al., 2023; Deng et al., 2024; Zhou et al., 2024b), software manipulation (Kapoor et al., 2024; Rawles et al., 2023; Yang et al., 2023b), and gaming (Ma et al., 2023b; Wang et al., 2023h; b; Xu et al., 2024) have emerged. In the gaming realm, agents have been deployed in simulated interactive environments, including Minecraft (Wang et al., 2023h; b), Starcraft II (Ma et al., 2023b), and Red Dead Redemption (Tan et al., 2024). These agents receive textual observations from internal APIs and execute predefined semantic actions. However, these domain-specific applications limit the agents’ ability to generalize to other games or broader software applications. While Wang et al. (2023a) attempts to interact with the environment in a human-like manner using screenshots as input and controlling mouse and keyboard actions, its action space is constrained to a predefined hybrid space. Despite achieving notable results in specific tasks, these methods lack the ability to generalize across diverse tasks due to inconsistencies in observation and action spaces. Several research efforts (Zhang et al., 2024b; Gao et al., 2023b; Kapoor et al., 2024; Niu et al., 2024; Wu et al., 2024b) have aimed to enhance the scalability of web agents by utilizing screenshots as input and keyboard and mouse operations as output, allowing them to interact with a wider range of applications.
@@
  • Advanced interface encodes unified representations. The integration of AGI interface is paramount for enhancing model performance. These interfaces can supervise less capable models in sequential interactions and introduce a range of novel capabilities in parallel cooperation. A key focus should be developing comprehensive and lightweight AGI interfaces that exhibit strong generative and understanding abilities. Future advanced interfaces should be able to encode a unified representation across all modalities. This would simplify the aligning process and optimize resource allocation for various downstream tasks, such as generation and identification.
@@
  • Advanced agent network promotes cooperative learning. Enhancing the network of AGI agents to foster social and cooperative learning is essential for the advancement of collective intelligence. By enabling agents to share insights, strategies, and knowledge, we can facilitate a more rapid and efficient learning process across different domains. Social learning mechanisms, such as imitation and observation, can be incorporated to enable agents to learn from the successes and failures of their peers. Furthermore, cooperative learning models can be designed to encourage agents to work together towards common goals, harnessing their diverse strengths and capabilities. Such a networked approach not only accelerates the pace of innovation but also leads to developing more versatile and adaptive AGI systems capable of tackling complex, real-world problems through teamwork and collaboration.
@@ 4 AGI Systems: Implementing the Mechanism of AGI
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=3.5cm, s sep=10pt, calign=child edge, calign child=(n_children()+1)/2, [AGI Systems, fill=red!20 [Scalable Model
@@ 4.1 System Challenges
Privacy-sensitive and Resource-constraint Settings While current most successful large models are deployed in data centers where requests from users are processed in a centralized manner (Achiam et al., 2023; AI, 2024), the need for serving models on edges where data and queries are used and processed locally in more privacy or latency-sensitive situations will become more substantial. However, edge devices are usually less capable in terms of computing and memory, which motivates developing techniques that optimize the utilization under resource-constraint settings and, at the same time, do not compromise model performance.
@@ 4.2 Scalable Model Architectures
Self-attention Patterns Vanilla self-attention, the core mechanism under transformer architectures (Vaswani et al., 2017), scales quadratically with the input sequence length, drastically limiting its potential on long context tasks with current hardware memory capacity. However, the full causal mask used in self-attention might not be efficiently optimal (Farina et al., 2024), wasting computation and time. Sharing similar insights to the lottery ticket hypothesis Frankle and Carbin (2019) and the structured sparsity observation (Dong et al., 2023a), many works explored different attention mask patterns to reduce the computations over the fully casual one. Sliding window Beltagy et al. (2020); Jiang et al. (2023d) and dilated patterns Ding et al. (2023b) either limit to local attention or reduce the resolution, often resulting in improving efficiency with some performance degradation. Another line of work observes that some tokens in the sentence are semantically more important and hence introduce different kinds of global tokens for efficient attention, such as the initial tokens (Xiao et al., 2023) and the special landmark (Mohtashami and Jaggi, 2023) for each block of tokens. It is worth noting that the computation bottleneck might switch between the self-attention and fully connected layers for models with different scales, and blindly applying heuristics-based sparse patterns might not only give marginal speedup but also incur a loss in performance, which motivates future research in more complex and adaptive efficient patterns.
@@
Model Compression The goal of model compression is to reduce the memory footprints, computational complexity, and deployment cost of large models. Among many approaches, knowledge distillation (KD) (Hinton et al., 2015; Hsieh et al., 2023) is the process of extracting the learned knowledge of a bigger teacher model to a smaller student model that can often be served more efficiently. On the one hand, black-box LLM knowledge distillation only requires querying the teacher model (via API calls) and collecting input and output pairs which can be used to train the student model. Following this line, Alpaca (Taori et al., 2023) only costs around 100 US dollars to balance efficiency and performance by distilling from ChatGPT. Vicuna (Peng et al., 2023b) shows promising performance on instruction fine-tuning with knowledge distillation from GPT-4. On the other hand, MiniLLM (Gu et al., 2023) explores white-box knowledge distillation where more information like model weights and loss values of the teacher model is accessible, which can potentially stimulate better knowledge transfer. MiniLLM proposes to replace the standard KL divergence objective with reverse KL which prevents the student model from overestimating the low-probability regions of the teacher distribution. Balancing the level of access to the teacher model will remain as a relevant topic for algorithmic design, safety, data privacy, cost, and efficiency. It is also exciting to explore the possibility of reverse-learning or super-alignment  [^2] https://openai.com/blog/introducing-superalignment where we want to distill knowledge from weaker models that can be leveraged (e.g., through analyzing, merging, and adaptively updating) to improve the current one.
@@
Beyond Transformers Despite the enormous success of the ubiquitous transformer architecture, many works attempted to find other designs to overcome some of its shortcomings. Mixture of Experts (MoEs) (Shazeer et al., 2017; Roy et al., 2020) replace the dense layers in transformer models with a conditional module consisting of multiple “expert” sub-networks. A routing mechanism is used to dynamically decide which expert(s) to use on the token-level (Zhou et al., 2022a; Fedus et al., 2022) or task-level (Kudugunta et al., 2021). Despite having multiple experts, sparse MoEs can often train and decode faster with the same model size and are expected to specialize in different abstract tasks (Jiang et al., 2023d; Hwang et al., 2023; Gale et al., 2022). However, MoEs also raise other system challenges during inference such as higher requirements for loading all experts into the VRAM and distribute experts over multiple nodes.
@@
State space models (SSMs) have recently been applied to model sequence-to-sequence transformations (Gu et al., 2022a; Gupta et al., 2022; Smith et al., 2023), which can be readily used in various model architecture topology to replace the quadratic self-attention mechanism. A (discretized) SSM defines a recurrence relationship along each time-step (token) via a tuple of learnable parameters $(\Delta,\bar{A},\bar{B},C)$ and the major challenge that most works try to solve is how to compute this recurrence in a parallelizable way that can efficiently use modern hardware accelerators (e.g. FFTConv). The simplest form in this category is linear attention (Katharopoulos et al., 2020; Yang et al., 2023e), which can be viewed as a degenerate SSM. In its core, linear attention expresses self-attention as a linear dot-product of kernel feature maps and make use of the associativity property of matrix products to reduce the complexity down to linear. S4 (Gu et al., 2022a) parameterizes the SSM both expressively and efficiently via a low-rank correction, allowing it to be diagonalized stably and reducing the SSM to the well-studied computation of a Cauchy kernel. There are many other following works (Gupta et al., 2022; Gu et al., 2022b; Smith et al., 2023) after S4 which attempt different parameterization of the transition matrix $\bar{A}$ (and others) to improve both the computational efficiency and modeling capacities. H3 (Fu et al., 2023a) proposes a SSM block that consists of two stacked separate SSMs that are specially designed to meet the challenge of recalling earlier tokens and support token comparisons across sequences. Hyena (Poli et al., 2023) generalizes H3 by replacing the S4 layer with an interleaved and implicitly parameterized long convolutions and data-controlled gating, which disentangles parameter size from the filter size and hence allows for greater expressivity. (Sun et al., 2023a) proposes Retentive Network a foundation architecture that includes additional gates and uses a variant of multi-head attention, achieving impressive constant inference cost and linear long-sequence memory consumption. RWKV (Peng et al., 2023a; 2024) is a new architecture that takes advantage of the efficient parallelizable training of Transformers with the efficient inference of RNNs. In essence, the main “WKV” operation involves linear time invariance (LTI) recurrences, which can be interpreted as a ratio of two SSMs (Gu and Dao, 2023). To tackle the key weakness of previous SSM models which is their inability to perform content-based reasoning, Mamba (Gu and Dao, 2023) proposes the selective state space that can make the SSM parameters as functions of the input, effectively turning the SSM from LTI to time-varying. Despite no longer being able to apply efficient convolutions, they designed a hardware-aware parallel algorithms for recurrence computation called Parallel Associative Scan, which enables it to achieve over 5×5$\times$5 × higher throughput than Transformers, SoTA performance across several modalities, and keep improving on real data up to million-length sequences.
@@
Recent works also explored hybrid architectures that wish to bring the benefits from different variants. (Lieber et al., 2024) proposes to combine Transformers with Mamba by interleaving layers, which achieves impressive results on both standard and long context tasks with manageable resource requirements. Beyond manual design, MAD (Poli et al., 2024) integrates the process into an end-to-end pipeline consisting of small-scale capability unit tests predictive of scaling laws. MAD successfully finds an efficient architecture, called Striped Hyena, based on hybridization and sparsity, which outperforms state-of-the-art Transformer, convolutional, and recurrent architectures (Transformer++, Hyena, Mamba) in scaling, both at compute-optimal budgets and in over-trained regimes. These works will likely continue to inspire further explorations in architectural designs that are both performant and efficient at scaling, breaking through the current Transformer paradigm.

### v2

@@ 3.3.1 AI Interface to Other AI agents
  • Agents as teachers and learners. On one hand, stronger AGI models often act as oracles to provide ‘supervision’ to inferior ones, from tuning on data from better models (Taori et al., 2023; Gu et al., 2023) to prompt-engineering-based approaches (Huang et al., 2022a; Jiang et al., 2023a; Fu et al., 2023b), there emerges the concept of model knowledge distillation. In the field of language processing, it is common to use a teacher system to label and expand existing data by directly taking the teacher’s answer (Gilardi et al., 2023; Hsieh et al., 2023; Li et al., 2022a; Sun et al., 2024b; Ding et al., 2023a) or using more advanced techniques such as CoT prompting (Ramnath et al., 2023; Li et al., 2023g), or creating new data for subsequent models to distill useful and compact knowledge from large-scale data (Li et al., 2023b; Javaheripi et al., 2023). Similar paradigms are applied in computer vision and multimodal domains for better model training and deployment. One of the most prevailing methods is utilizing the GPT-4V model to label answers in various tasks (Shu et al., 2023; Liu et al., 2023d; Li et al., 2023j).
    In conventional solutions, a better AGI model severs as the role of the teacher, however, there is a growing trend for less competent AGI models to provide insights for aligning stronger ones with or beyond human perception, known as superalignment (Burns et al., 2023). This has proven to be effective in empowering higher capacity than the teacher model with a vanilla fine-tuning strategy and distilled data from the smaller model. Recent studies have begun to explore how leveraging weaker models to enhance the performance of stronger ones can be applied across various domains. Chen et al. (2024a) introduced the innovative idea of iteratively harnessing the capacity of weaker models to improve the efficacy of more powerful counterparts. (Ji et al., 2024) developed an efficient alignment paradigm that learns the correctional residuals between aligned and unaligned responses from a weaker model. (Sun et al., 2024c) employs a less advanced system, trained on simpler tasks, to guide more capable models for tackling hard reasoning tasks (e.g., level 4-5 MATH problems). In the vision domain, (Guo et al., 2024a) presented compelling evidence that weak-to-strong generalization may outperform finetuning methods in certain scenarios. This interaction between different models requires acquiring knowledge from selected models step-by-step, forming the sequential interface.
@@
    Looking beyond multi-modal tasks, more specific and advanced agent systems that can handle more complex computer applications such as web browsing (Mialon et al., 2023; Deng et al., 2024; Zhou et al., 2024b), software manipulation (Kapoor et al., 2024; Rawles et al., 2023; Yang et al., 2023b), and gaming (Ma et al., 2023b; Wang et al., 2023h; b; Xu et al., 2024) have emerged. In the gaming realm, agents have been deployed in simulated interactive environments, including Minecraft (Wang et al., 2023h; b), Starcraft II (Ma et al., 2023b). These agents receive textual observations from internal APIs and execute predefined semantic actions. However, these domain-specific applications limit the agents’ ability to generalize to other games or broader software applications. Tan et al. (2024) proposes the concept of General Computer Control (GCC) and adopts a more intuitive approach, employing multimodal input from screenshots to generate keyboard and mouse commands within Red Dead Redemption II. This setting holds promise for expansion into more intricate computer tasks. While Wang et al. (2023a) attempts to interact with the environment in a human-like manner using screenshots as input and controlling mouse and keyboard actions, its action space is constrained to a predefined hybrid space. Despite achieving notable results in specific tasks, these methods lack the ability to generalize across diverse tasks due to inconsistencies in observation and action spaces. Several research efforts (Zhang et al., 2024b; Gao et al., 2023b; Kapoor et al., 2024; Niu et al., 2024; Wu et al., 2024b) have aimed to enhance the scalability of web agents by utilizing screenshots as input and keyboard and mouse operations as output, allowing them to interact with a wider range of applications.
@@
  • Advanced interface encodes unified representations. The integration of the AGI interface is paramount for enhancing model performance. These interfaces can supervise less capable models in sequential interactions and introduce a range of novel capabilities in parallel cooperation. A key focus should be developing comprehensive and lightweight AGI interfaces that exhibit strong generative and understanding abilities. Future advanced interfaces should be able to encode a unified representation across all modalities. This would simplify the aligning process and optimize resource allocation for various downstream tasks, such as generation and identification.
@@
  • Advanced agent network promotes cooperative learning. Enhancing the network of AGI agents to foster social and cooperative learning is essential for the advancement of collective intelligence. By enabling agents to share insights, strategies, and knowledge, we can facilitate a more rapid and efficient learning process across different domains. Social learning mechanisms, such as imitation and observation, can be incorporated to enable agents to learn from the successes and failures of their peers. Furthermore, cooperative learning models can be designed to encourage agents to work together towards common goals, harnessing their diverse strengths and capabilities. Such a networked approach not only accelerates the pace of innovation but also leads to the developing more versatile and adaptive AGI systems capable of tackling complex, real-world problems through teamwork and collaboration.
@@ 4 AGI Systems: Implementing the Mechanism of AGI
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=3.5cm, s sep=10pt, calign=child edge, [AGI Systems, fill=red!20 [Scalable Model
@@ 4.1 System Challenges
Privacy-sensitive and Resource-constraint Settings While the current most successful large models are deployed in data centers where requests from users are processed in a centralized manner (Achiam et al., 2023; AI, 2024), the need for serving models on edges where data and queries are used and processed locally in more privacy or latency-sensitive situations will become more substantial. However, edge devices are usually less capable in terms of computing and memory, which motivates developing techniques that optimize the utilization under resource-constraint settings and, at the same time, do not compromise model performance.
@@ 4.2 Scalable Model Architectures
Self-attention Patterns Vanilla self-attention, the core mechanism under transformer architectures (Vaswani et al., 2017), scales quadratically with the input sequence length, drastically limiting its potential on long context tasks with current hardware memory capacity. However, the full causal mask used in self-attention might not be efficiently optimal (Farina et al., 2024), wasting computation and time. The lottery ticket hypothesis Frankle and Carbin (2019) and the structured sparsity observation (Dong et al., 2023a) both suggest potential structural and computational redundancy in LLMs. Inspired by these works, many works explored different attention mask patterns to reduce the computations and memory requirement over the fully casual one. Sliding window Beltagy et al. (2020); Jiang et al. (2023d) and dilated patterns Ding et al. (2023b) either limit to local attention or reduce the resolution, often resulting in improving efficiency with some performance degradation. Another line of work observes that some tokens in the sentence are semantically more important and hence introduce different kinds of global tokens for efficient attention, such as the initial tokens (Xiao et al., 2023) and the special landmark (Mohtashami and Jaggi, 2023) for each block of tokens. It is worth noting that the computation bottleneck might switch between the self-attention and fully connected layers for models with different scales, and blindly applying heuristics-based sparse patterns might not only give marginal speedup but also incur a loss in performance, which motivates future research in more complex and adaptive efficient patterns.
@@
Model Compression The goal of model compression is to reduce the memory footprints, computational complexity, and deployment cost of large models. Among many approaches, Knowledge Distillation (KD) (Hinton et al., 2015; Hsieh et al., 2023) is the process of extracting the learned knowledge of a bigger teacher model to a smaller student model that can often be served more efficiently. On the one hand, black-box LLM knowledge distillation only requires querying the teacher model (via API calls) and collecting input and output pairs which can be used to train the student model. Following this line, Alpaca (Taori et al., 2023) only costs around 100 US dollars to balance efficiency and performance by distilling from ChatGPT. Vicuna (Peng et al., 2023b) shows promising performance on instruction fine-tuning with knowledge distillation from GPT-4. On the other hand, MiniLLM (Gu et al., 2023) explores white-box knowledge distillation where more information like model weights and loss values of the teacher model is accessible, which can potentially stimulate better knowledge transfer. MiniLLM proposes to replace the standard KL divergence objective with reverse KL which prevents the student model from overestimating the low-probability regions of the teacher distribution. Generalized Knowledge Distillation (GKD) (Agarwal et al., 2024) breaks the traditional supervised KD training region by taking advantage of student’s self-generated output sequences and leveraging feedback from the teacher on such sequences. This not only helps mitigate the distribution mismatch between training and inference but also has been proven to be more useful when the student model lacks enough capacity to fully mimic the teacher’s behaviors. Balancing the level of access to the teacher model will be remained as a relevant topic for algorithmic design, safety, data privacy, cost, and efficiency. It is also exciting to explore the possibility of reverse-learning or super-alignment  [^2] https://openai.com/blog/introducing-superalignment where we want to distill knowledge from weaker models that can be leveraged (e.g., through analyzing, merging, and adaptively updating) to improve the current one.
@@
Beyond Transformers Despite the enormous success of the ubiquitous transformer architecture, many works attempted to find other designs to overcome some of its shortcomings. Mixture of Experts (MoEs) (Shazeer et al., 2017; Roy et al., 2020) replace the dense layers in transformer models with a conditional module consisting of multiple “expert” sub-networks. A routing mechanism is used to dynamically decide which expert(s) to use on the token-level (Zhou et al., 2022a; Fedus et al., 2022) or task-level (Kudugunta et al., 2021). Despite having multiple experts, sparse MoEs can often train and decode faster with the same model size and are expected to specialize in different abstract tasks (Jiang et al., 2023d; Hwang et al., 2023; Gale et al., 2022). However, MoEs also raise other system challenges during inference such as higher requirements for loading all experts into the VRAM and distributing experts over multiple nodes.
@@
State space models (SSMs) have recently been applied to model sequence-to-sequence transformations (Gu et al., 2022a; Gupta et al., 2022; Smith et al., 2023), which can be readily used in various model architecture topology to replace the quadratic self-attention mechanism. A (discretized) SSM defines a recurrence relationship along each time-step (token) via a tuple of learnable parameters $(\Delta,\bar{A},\bar{B},C)$ and the major challenge that most works try to solve is how to compute this recurrence in a parallelizable way that can efficiently use modern hardware accelerators (e.g. FFTConv). The simplest form in this category is linear attention (Katharopoulos et al., 2020; Yang et al., 2023e), which can be viewed as a degenerate SSM. At its core, linear attention expresses self-attention as a linear dot-product of kernel feature maps and makes use of the associativity property of matrix products to reduce the complexity down to linear. S4 (Gu et al., 2022a) parameterizes the SSM both expressively and efficiently via a low-rank correction, allowing it to be diagonalized stably and reducing the SSM to the well-studied computation of a Cauchy kernel. There are many other following works (Gupta et al., 2022; Gu et al., 2022b; Smith et al., 2023) after S4 which attempt different parameterization of the transition matrix $\bar{A}$ (and others) to improve both the computational efficiency and modeling capacities. H3 (Fu et al., 2023a) proposes an SSM block that consists of two stacked separate SSMs that are specially designed to meet the challenge of recalling earlier tokens and support token comparisons across sequences. Hyena (Poli et al., 2023) generalizes H3 by replacing the S4 layer with an interleaved and implicitly parameterized long convolutions and data-controlled gating, which disentangles parameter size from the filter size and hence allows for greater expressivity. (Sun et al., 2023a) proposes Retentive Network a foundation architecture that includes additional gates and uses a variant of multi-head attention, achieving impressive constant inference cost and linear long-sequence memory consumption. RWKV (Peng et al., 2023a; 2024a) is a new architecture that takes advantage of the efficient parallelizable training of Transformers with the efficient inference of RNNs. In essence, the main “WKV” operation involves linear time invariance (LTI) recurrences, which can be interpreted as a ratio of two SSMs (Gu and Dao, 2023). To tackle the key weakness of previous SSM models which is their inability to perform content-based reasoning, Mamba (Gu and Dao, 2023) proposes the selective state space that can make the SSM parameters as functions of the input, effectively turning the SSM from LTI to time-varying. Despite no longer being able to apply efficient convolutions, they designed hardware-aware parallel algorithms for recurrence computation called Parallel Associative Scan, which enables it to achieve over 5×5$\times$5 × higher throughput than Transformers, SoTA performance across several modalities, and keep improving on real data up to million-length sequences.
Sparkles of interest for revisiting recurrent neural networks (RNNs) have also emerged with its major advantage in long context processing (i.e. linear time and constant memory for the hidden state). One of the challenges for RNNs is to scale the training and inference efficiently. (De et al., 2024) introduces Hawk, an RNN-based model with gated linear recurrences, and Griffin which mixes gated linear recurrences with local attention. They showcase the superior performance of Hawk against Mamba on downstream tasks, and Griffin matches performance of Llama-2 with six times fewer training tokens. Not only do they demonstrate the potential of long context capability, but also explain how to effectively leverage hardware accelerators during distributed training and inference by scaling Griffin to 14B parameters. Right after that, a family of models called RecurrentGemma (Botev et al., 2024) came out with various model sizes, in both pretrained and instruct-tuned versions. These advancements present the possibility of training a data-efficient, fixed state size, long context, and expressive model without relying solely on transformer architectures.
@@
Recent works also explored high-level architecture hybridization strategies that wish to bring the benefits from different variants. (Lieber et al., 2024) proposes to combine Transformers with Mamba by interleaving layers, which achieves impressive results on both standard and long context tasks with manageable resource requirements. Beyond manual design, MAD (Poli et al., 2024) integrates the process into an end-to-end pipeline consisting of small-scale capability unit tests predictive of scaling laws. MAD successfully finds an efficient architecture, called Striped Hyena, based on hybridization and sparsity, which outperforms state-of-the-art Transformer, convolutional, and recurrent architectures (Transformer++, Hyena, Mamba) in scaling, both at compute-optimal budgets and in over-trained regimes. These works will likely continue to inspire further explorations in architectural designs that are both performant and efficient at scaling, breaking through the current Transformer paradigm.